5a = 2b, 2a + 11 = 3b
Answer by ptth176(50)

5a = 2b=>a=2/5 b
2a + 11 = 3b=>2(2/5)b+11=3b
=>4/5b +11=3b
=>11/5b=11
=>b=5
=>a(2/5)5=2